# Testing quantum gravity effects with global lepton number symmetry and dark matter

### v1

@@ V.1 Analysis with dimension 7 operator
We choose a different benchmark point (BP3) to illustrate the combined result for DM and GW phenomenology. The chosen values for BP3 are : $v_{\Phi}=10^{12}$ GeV, $\lambda_{\Phi}=10^{-9}$, $m_{N}=200$ GeV, $m_{\chi}=100$ MeV, $\Lambda_{\rm QG}=M_{\rm Pl}$ and majoron mass in sub-eV scale. For BP3 the mass of $\phi$ is found to be $m_{\phi}=4.5\times 10^{7}$ GeV. Although to satisfy correct DM relic, it is sufficient to consider $2m_{N}<m_{\phi}$ as discussed in sub-section III.1 and III.2, $m_{N}$ around TeV scale requires a DM mass around GeV scale (e.g., $m_{N}=10^{4}$ GeV requires $m_{\chi}\sim 1$ GeV). Now, a GeV scale DM mass along with $v_{\Phi}=10^{12}$ GeV and $\Lambda_{\rm QG}=M_{\rm Pl}$ results in short DM lifetime in tension with indirect detection limits. Therefore, we choose a smaller $m_{N}=200$ GeV and hence a smaller $m_{\chi}=100$ MeV for BP3. The evolution of relevant number densities together with DM number density for BP3 are shown in the left plot of Fig. 8. The final freeze-in abundance of DM, dominantly produced from decay, is consistent with the observed DM abundance. In the right panel of Fig. 8, we summarise our result for the dimension 7 operator. Here, we notice that the QG scale can be reduced to the Planck scale but only for a relatively large value of $v_{\Phi}$ and a smaller value of $\lambda_{\Phi}$. Note that the price we pay to achieve the desired DM and GW phenomenology is a relatively small RHN mass (sub-TeV). As a result, its interaction with majoron becomes very feeble and the majorons never thermalises. This is the reason why we do not see the $\Delta N_{\rm eff}$ bounds from thermalised majoron in Fig. 8. Despite this, the parameter space and BP3 in particular, remain within reach of near-future GW experiments like LISA and DM indirect detection experiments.
@@
Figure 8: Left panel: evolution of comoving number densities for BP3: $v_{\Phi}=10^{12}$ GeV, $\lambda_{\Phi}=10^{-9}$ (which gives $m_{\phi}=4.5\times 10^{7}$ GeV), $m_{N}=200$ GeV and $m_{\chi}=100$ MeV. Right panel: combined constraints on $\Lambda_{\rm QG}$ and $v_{\Phi}$ with $\lambda_{\Phi}=10^{-9}$. The light gray and yellow-shaded regions denote the excluded regions by $\gamma$-ray and CMB observations respectively with DM mass 100 MeV. The black dashed line shows the sensitivity of upcoming SKA telescope. The black dot corresponds to BP3. The red, blue, brown, cyan, purple, and green curves represent the sensitivities of THEIA, SKA, DECIGO, LISA, ET and IPTA with SNR = 10. The gray-shaded regions are excluded by the requirement of BBN. The pink-shaded region corresponds to the scenario where DW may overclose the Universe at an early epoch.
@@ VI Conclusion
We have studied the possibility of testing quantum gravity effects in the singlet majoron model with dark matter. The spontaneous breaking of global lepton number symmetry generates right-handed neutrino mass or type-I seesaw scale dynamically. Explicit $U(1)_{L}$ breaking of the quadratic type not only leads to the mass of majoron, the pseudo-NG boson but also reduces the $U(1)_{L}$ symmetry to a $Z_{2}$. The effective spontaneous breaking of $Z_{2}$ symmetry then leads to the formation of domain walls. We consider higher dimensional $Z_{2}$-breaking operators suppressed by the QG scale as bias necessary for the domain walls to annihilate away while generating stochastic gravitational waves. The same QG scale-suppressed operators also lead to the decay of dark matter. In the minimal singlet majoron model, such QG scale suppressed operators can arise at dimension five level and get tightly constrained from DW, GW as well as DM lifetime criteria. These constraints make the non-thermal production of DM via singlet scalar portal couplings more realistic. We consider the majoron to be in the sub-eV regime and produced thermally. This brings another detection prospect as light majoron can lead to observable $\Delta N_{\rm eff}$ within reach of future CMB experiments. While the scale of QG is required to be $\Lambda_{\rm QG}\gtrsim 10^{25}$ GeV for detectable GW from DW annihilation while satisfying indirect detection bounds on decaying DM, it can be brought down to the Planck scale if explicit global symmetry operators are considered at dimension 7 level instead of dimension 5.
@@ Appendix B Cross-section
$-\frac{2\left(-16m_{N}^{4}+4m_{N}^{2}s+m_{\phi}^{4}\right)\log\left(\frac{s\left(\sqrt{\frac{\left(s-4m_{N}^{2}\right)\left(s-4m_{\phi}^{2}\right)}{s^{2}}}-1\right)+2m_{\phi}^{2}}{2m_{\phi}^{2}-s\left(\sqrt{\frac{\left(s-4m_{N}^{2}\right)\left(s-4m_{\phi}^{2}\right)}{s^{2}}}+1\right)}\right)}{2m_{\phi}^{2}-s}\Biggl{)}$ (34)

### v2

@@ V.1 Analysis with dimension 7 operator
We choose a different benchmark point (BP3) to illustrate the combined result for DM and GW phenomenology. The chosen values for BP3 are : $v_{\Phi}=10^{12}$ GeV, $\lambda_{\Phi}=10^{-9}$, $m_{N}=20$ GeV, $m_{\chi}=20$ MeV, $\Lambda_{\rm QG}=M_{\rm Pl}$ and majoron mass in sub-eV scale. For BP3 the mass of $\phi$ is found to be $m_{\phi}=4.5\times 10^{7}$ GeV. Although to satisfy correct DM relic, it is sufficient to consider $2m_{N}<m_{\phi}$ as discussed in sub-section III.1 and III.2, $m_{N}$ around TeV scale requires a DM mass around GeV scale (e.g., $m_{N}=10^{4}$ GeV requires $m_{\chi}\sim 1$ GeV). Now, a GeV scale DM mass along with $v_{\Phi}=10^{12}$ GeV and $\Lambda_{\rm QG}=M_{\rm Pl}$ results in short DM lifetime in tension with indirect detection limits. Therefore, we choose a smaller $m_{N}=20$ GeV and hence a smaller $m_{\chi}=20$ MeV for BP3. The evolution of relevant number densities together with DM number density for BP3 are shown in the left plot of Fig. 8. The final freeze-in abundance of DM, dominantly produced from decay, is consistent with the observed DM abundance. In the right panel of Fig. 8, we summarise our result for the dimension 7 operator. Here, we notice that the QG scale can be reduced to the Planck scale but only for a relatively large value of $v_{\Phi}$ and a smaller value of $\lambda_{\Phi}$. Note that the price we pay to achieve the desired DM and GW phenomenology is a relatively small RHN mass (sub-TeV). As a result, its interaction with majoron becomes very feeble and the majorons never thermalises. This is the reason why we do not see the $\Delta N_{\rm eff}$ bounds from thermalised majoron in Fig. 8. Despite this, the parameter space and BP3 in particular, remain within reach of near-future GW experiments like LISA and DM indirect detection experiments. While light majoron does not give rise to observable $\Delta N_{\rm eff}$ in this scenario, higher dimensional operators can make majoron very heavy due to lower values of QG scale. If majoron becomes heavier than DM, we can have additional freeze-in contribution to DM relic. We find that this contribution is of same order as the contribution from the decay of $\phi$ without changing the generic conclusions. The details of majoron decay contribution to DM relic is given in appendix C.
@@
Figure 8: Left panel: evolution of comoving number densities for BP3: $v_{\Phi}=10^{12}$ GeV, $\lambda_{\Phi}=10^{-9}$ (which gives $m_{\phi}=4.5\times 10^{7}$ GeV), $m_{N}=20$ GeV and $m_{\chi}=20$ MeV. Right panel: combined constraints on $\Lambda_{\rm QG}$ and $v_{\Phi}$ with $\lambda_{\Phi}=10^{-9}$. The light gray and yellow-shaded regions denote the excluded regions by X-ray and CMB observations respectively with DM mass 20 MeV. The black dashed line shows the sensitivity of upcoming SKA telescope. The black dot corresponds to BP3. The red, blue, brown, cyan, purple, and green curves represent the sensitivities of THEIA, SKA, DECIGO, LISA, ET and IPTA with SNR = 10. The gray-shaded regions are excluded by the requirement of BBN. The pink-shaded region corresponds to the scenario where DW may overclose the Universe at an early epoch.
@@ VI Conclusion
We have studied the phenomenology of a dark matter scenario embedded in a global lepton number symmetry $U(1)_{L}$ focusing on the detection prospects. The spontaneous breaking of global lepton number symmetry generates right-handed neutrino mass or type-I seesaw scale dynamically. Explicit $U(1)_{L}$ breaking of the quadratic type not only leads to the mass of majoron, the pseudo-NG boson but also reduces the $U(1)_{L}$ symmetry to a $Z_{2}$. The effective spontaneous breaking of $Z_{2}$ symmetry then leads to the formation of domain walls. We consider higher dimensional $Z_{2}$-breaking operators suppressed by the QG scale as bias necessary for the domain walls to annihilate away while generating stochastic gravitational waves. The same QG scale-suppressed operators also lead to the decay of dark matter. In the minimal singlet majoron model, such QG scale suppressed operators can arise at dimension five level and get tightly constrained from DW, GW as well as DM lifetime criteria. These constraints make the non-thermal production of DM via singlet scalar portal couplings more realistic. We consider the majoron to be in the sub-eV regime and produced thermally. This brings another detection prospect as light majoron can lead to observable $\Delta N_{\rm eff}$ within reach of future CMB experiments. While the scale of QG is required to be $\Lambda_{\rm QG}\gtrsim 10^{25}$ GeV for detectable GW from DW annihilation while satisfying indirect detection bounds on decaying DM, it can be brought down to the Planck scale if explicit global symmetry operators are considered at dimension 7 level instead of dimension 5.
@@ Appendix B Cross-section
Here we provide the expressions for relevant cross-sections used in the numerical analysis.
@@
$-\frac{2\left(-16m_{N}^{4}+4m_{N}^{2}s+m_{\phi}^{4}\right)\log\left(\frac{s\left(\sqrt{\frac{\left(s-4m_{N}^{2}\right)\left(s-4m_{\phi}^{2}\right)}{s^{2}}}-1\right)+2m_{\phi}^{2}}{2m_{\phi}^{2}-s\left(\sqrt{\frac{\left(s-4m_{N}^{2}\right)\left(s-4m_{\phi}^{2}\right)}{s^{2}}}+1\right)}\right)}{2m_{\phi}^{2}-s}\Biggl{)}.$ (35)
Appendix C Dark matter from Majoron decay
In this work, we have focused on light majorons which remain relativistic during recombination and contribute to $\Delta N_{\rm eff}$. Therefore, we have kept the soft $U(1)_{L}$ breaking mass term in the sub-eV range. However, higher dimensional operators suppressed by powers of $\Lambda_{\rm QG}$ can push the majoron to higher masses. In the minimal model with dimension five operators suppressed by $\Lambda_{\rm QG}$, the next-to-leading-order (NLO) contribution to majoron mass is $\sim\sqrt{v^{3}_{\Phi}/\Lambda_{\rm QG}}$. For some parts of the parameter space, this contribution can be kept at or below eV scale while for the rest of the parameter space, some tuning will be required between bare mass and NLO mass such that the net majoron mass remains at or below the eV scale. This keeps the $\Delta N_{\rm eff}$ prospects alive as discussed above.
On the other hand, for lower values of $\Lambda_{\rm QG}$ as applicable in a scenario where NLO terms arise at dimension seven order, majoron mass $\sim\sqrt{v^{5}_{\Phi}/\Lambda^{3}_{\rm QG}}$ can become substantially large. Since we do not have observable $\Delta N_{\rm eff}$ in the scenario with dimension seven operators as discussed earlier, we do not lose any interesting phenomenology if majorons become heavy and non-relativistic due to NLO operators. However, if majorons become even heavier than DM that is, $m_{\eta}>2m_{\chi}$, majoron can decay to DM and contribute to the DM abundance. The coupled BEs for majoron and DM can be written as
$\frac{dY_{\eta}}{dx}$ $=$ $-\frac{\beta s}{\mathcal{H}x}\langle\sigma v\rangle_{\eta\eta\to N\bar{N}}(Y^{2}_{\eta}-(Y^{\rm eq}_{\eta})^{2})-\frac{\beta}{\mathcal{H}x}\Gamma_{\eta\to N\bar{N}}\frac{K_{1}(m_{\eta}/T)}{K_{2}(m_{\eta}/T)}(Y_{\eta}-Y^{\rm eq}_{\eta})$
$-\frac{\beta}{\mathcal{H}x}\Gamma_{\eta\to\chi\bar{\chi}}\frac{K_{1}(m_{\eta}/T)}{K_{2}(m_{\eta}/T)}Y_{\eta},$
$\frac{dY_{\chi}}{dx}$ $=$ $\frac{\beta}{\mathcal{H}x}\Gamma_{\phi\to\chi\bar{\chi}}\frac{K_{1}(m_{\phi}/T)}{K_{2}(m_{\phi}/T)}Y_{\phi}+\frac{\beta}{\mathcal{H}x}\Gamma_{\eta\to\chi\bar{\chi}}\frac{K_{1}(m_{\eta}/T)}{K_{2}(m_{\eta}/T)}Y_{\eta}+\frac{\beta s}{\mathcal{H}x}\langle\sigma v\rangle_{N\bar{N}\to\chi\bar{\chi}}(Y^{\rm eq}_{N})^{2}.$ (37)
To compare the production of DM from majoron decay and from $\phi$ decay, we plot the comoving number densities for BP3 in Fig. 9. For BP3, $v_{\Phi}=10^{12}$ GeV, $\lambda_{\Phi}=10^{-9}$ ($m_{\phi}=4.5\times 10^{7}$ GeV), $m_{N}=20$ GeV, $m_{\chi}=20$ MeV, $\Lambda_{\rm QG}=M_{\rm Pl}$. The majoron mass is determined by the dimension $7$ operator which is calculated as $m_{\eta}=\sqrt{\frac{21}{4\sqrt{2}}\frac{v_{\Phi}^{5}}{\Lambda_{\rm QG}^{3}}}\simeq 61\text{ GeV}$. Fig. 9 shows that the DM yield from $\phi$ and majoron $\eta$ decay by solid blue and dashed blue lines respectively. About equal amount of DM is produced from both the $\phi$ and majoron decay as the the branching ratio of $\phi\to\chi\bar{\chi}$ and $\eta\to\chi\bar{\chi}$ are similar. However, the production of DM from majoron decay is delayed compared to $\phi$ decay due to smaller decay width of majoron.
Figure 9: Evolution of comoving number densities of different species for BP3. The solid red and solid green color lines denote the evolution of $\phi$ and $\eta$ respectively. DM abundance from $\phi$ and majoron decay are denoted by solid blue and dashed blue color lines respectively. Both $\phi$ and majoron decay contribute at the same order to the DM abundances as depicted in the figure.